OpenStudy (anonymous):

Find the value that makes the ratios equals. 36 to 1, ? to 1 (i need u to just check)

OpenStudy (anonymous):

i got 3 is that right

OpenStudy (anonymous):

are you saying: 36 to 1 is equal to ? to 1

OpenStudy (anonymous):

or did you make an error in your question? If you got 3, something is mixed up

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i ment 36 to 1 ? to 1

OpenStudy (anonymous):

Remember, use the fraction form,\[\frac{ 36 }{ 1 } = 36\]Solve for x on the other one.

OpenStudy (anonymous):

\[\frac{ 36 }{ 1 } = \frac{ 36 }{ ? }\]

OpenStudy (anonymous):

i did and i got 3 so is that right

OpenStudy (anonymous):

no, read what jake wrote again~

OpenStudy (anonymous):

No be careful. If you input 3 for ?, then the results are not the same. You need a number in '?' which would equal with the left side.

OpenStudy (anonymous):

@partyrainbow276 you might be "over-thinking" this... look at the ratios above expressed as fractions... you can almost "see" what the "?" should be...

OpenStudy (anonymous):

is it 36

OpenStudy (anonymous):

@gezimbasha

OpenStudy (anonymous):

thx evry1